# Gravitational and electromagnetic Cherenkov radiation constraints in modified dispersion relations

### v1

@@ 4.2 Observational constraints in the gravitational sector
In order to obtain realistic bounds Kosteleck_2015 , let us consider that the UHECR is a proton of energy $p\sim 10^{11}\,\text{GeV}$. The source of these particles is still unknown Abreu_2022 but it is expected that are produced in active galactic nuclei; the nearest is found at a few Mpc, so we may take $ct\sim 10\,\text{Mpc}$. Here we do not take into account the spin of the proton and neither its inner structure. Regarding the former, we already have mentioned that spin corrections are not expected to affect the order of magnitude of the constraints on $\mathbb{A}_{\text{proton}}\equiv\mathbb{A}$ and $\mathbb{B}_{\text{graviton}}\equiv\mathbb{B}$. Attending to the latter, a detailed analysis of the gravitational CR emitted by a proton would require to consider its partonic structure. For soft emitted gravitons with $k\to 0$, this inner structure might be neglected and the pointlike approximation is valid Jacobson_2006 ; for hard emitted gravitons with $k\lesssim p$ it becomes important and the emission rate of its constituents should be computed. This would substantially increase the number of LIVPs considered as $\mathbb{A}_{\text{quark}}\neq\mathbb{A}_{\text{gluon}}$ in general, making the analysis more involved. This study is out of the scope of this work, but it is worth mentioning that this has been studied with more simple MDRs which consider $\alpha=\beta=0$ Gagnon_2004 ; Elliott_2005 . In short, we will assume that the proton has its own effective MDR, whose structure could be understood in terms of the composition of the MDRs of its constituents.
Computing the threshold momentum $k_{\text{th}}$ numerically for $m=m_{\text{proton}}$, performing the integral in equation (38) and imposing the condition $\text{d}E/\text{d}t\ll p/t$ we have obtained the allowed $\mathbb{A}-\mathbb{B}$ parameter space for $\alpha=2$, $\beta=0$ and $\alpha=1$, $\beta=2$ shown in figure 2. Values of $\mathbb{A}$ and $\mathbb{B}$ that define the dark-grey region permit Cherenkov emission without significant damping, whereas in the light-grey region the process is kinematically allowed but forbidden by the observation because of the damping effect. In both cases we see that no constraints are imposed in the region with $\mathbb{A}>0$ and $\mathbb{B}<0$, and also when $\mathbb{A},\mathbb{B}>0$ and $\mathbb{B}p^{\beta}-\mathbb{A}p^{\alpha}<0$. The latter is due to equation (18), since the threshold momentum is $k_{\text{min}}>p$. When both $\mathbb{A}$ and $\mathbb{B}$ are negative the region where kinematics does no permit Cherenkov emission is different in both cases. Attending to equation (17), the case where $\alpha=2$ and $\beta=0$ establishes that $\mathbb{B}-3\mathbb{A}p^{2}\geq 0$ so that the threshold momentum is $k_{\text{max}}>0$; by contrast, when $\alpha=1$ and $\beta=2$ the maximum momentum exists when $\mathbb{A}\leq 0$.
@@
The estimated excluded region appears in both cases for $|\mathbb{A}p^{\alpha}|\sim|\mathbb{B}p^{\beta}|\lesssim 10^{-17}$ when $p\sim 10^{11}\,\text{GeV}$ and $ct\sim 10\,\text{Mpc}$. Other positive integer values of $\alpha$ and $\beta$ exhibit constraints of the same order of magnitude for these energies and travelled distances. This allows us to check that LIVPs significantly modify the threshold condition of the process since $|\mathbb{A}p^{\alpha}|\gg(m/p)^{2}$, see equation (6). Hence, the approximations performed in section 2 are consistent and the analyses in the massless limit are accurate.
As we have previously advanced, figure 2 shows that it is possible to have arbitrarily large values of $\mathbb{A}$ and $\mathbb{B}$ where vacuum gravitational CR is allowed by observation as long as its difference is small enough. We have checked the same happens for many other positive integer values of $\alpha$ and $\beta$. This fact can be understood in the following way. When computing the energy loss rate, equation (38), it is necessary to obtain the integration limits from the threshold condition $\Theta_{c}=0$ for some given $\mathbb{A}$ and $\mathbb{B}$. Considering values of $\mathbb{A}$ and $\mathbb{B}$ where the process is allowed for some $k$, an increase in $\mathbb{A}$ and $\mathbb{B}$ will cause an increase in the value of $\Theta_{c}$ and thus in $\text{d}E/\text{d}t$. However, if the difference between $\mathbb{A}$ and $\mathbb{B}$ is small enough, the region of $k$ where the Cherenkov emission is possible will be narrower; this reduces the value of $\text{d}E/\text{d}t$ as the integration interval is smaller, and it can be such that the condition $\text{d}E/\text{d}t\ll p/t$ is still satisfied. Remember that $\Theta_{c}$ was obtained assuming $|\mathbb{A}p^{\alpha}|\sim|\mathbb{B}k^{\beta}|\ll 1$, and thus our model is not predictive for large values of $\mathbb{A},\mathbb{B}$.
@@ 4.3 Travel distance for particles emitting electromagnetic Cherenkov radiation
We have previously discussed that for positive integer values of $\alpha=\beta$ the energy loss rate of the particle $a$ due to vacuum electromagnetic CR is large enough so that the travelled distance for energies above the threshold $p_{\text{th}}$ is negligible compared to astrophysical distances Coleman_1997 ; Coleman_1999 ; Jacobson_2003 ; Jacobson_2006 . As our MDR does not assume equal values of the exponents of the particle $a$ and the photon ($\alpha$ and $\beta_{\gamma}$, respectively), it could be possible for $a$ to emit CR over large distances. Here we show that, even when $\alpha\neq\beta$, the vacuum electromagnetic CR is much more efficient than its gravitational counterpart, and that the particle $a$ losses all of its energy almost immediately above the threshold.
Figure 3: Constraints on the model for $\alpha=2$, $\beta=0$ (left panel) and $\alpha=1$, $\beta_{\gamma}=2$ (right panel) using an UHECR with $p\sim 10^{11}\,\text{GeV}$ and $ct\sim 0.1\,\text{fm}$. White, light-gray and dark-gray regions represent the same as in figure 2 but for the existence or absence of vacuum electromagnetic CR.
To illustrate this, let us consider an ultra high energy proton of energy $p\sim 10^{11}\,\text{GeV}$. As we do not solve the threshold equation (14) to find the threshold momentum $p_{\text{th}}$ of the particle $a$, we estimate that the order of magnitude of $\mathbb{A}$ and $\mathbb{B}_{\gamma}$ given by the kinematical constraints is the same as the one obtained in the gravitational case, see figure 2. For the same values of the exponents $\alpha$ and $\beta_{\gamma}$, figure 3 shows the parameter space $\mathbb{A}-\mathbb{B}_{\gamma}$ assuming a travelled distance of $ct\sim 0.1\,\text{fm}$. We see that the region where the damping is not significant lies in distance scales $39$ orders of magnitude smaller compared to those assumed for the emission of vacuum gravitational CR. This means that for UHECRs whose origin is galactic or extragalactic, the emission of electromagnetic CR occurs almost immediately after being produced, losing energy until reaching the threshold momentum $p_{\text{th}}$: the dark-grey region occupies an insignificant amount of the parameter space, and thus the kinematical analysis is enough to obtain bounds on the LIVPs. The estimation was made for two specific pair of values $(\alpha,\beta)$ but we have checked explicitly that similar results are obtained for many other positive integer values.
@@ 5 Conclusions
We would like to remark an important feature: the parameter space of the LI violating terms is not bounded, which implies that gravitational vacuum Cherenkov radiation is permitted as long as the difference between them is small enough. Different processes, such as the decay of a graviton in two high energy particles, could be used to restrict the allowed parameter space, but the quantum nature of gravity has not been resolved yet and no observations of gravitons have been made. Note that this problem is not present in the QED sector. On the one hand, the characteristic time of the analog processes mentioned is much smaller than the travelled distance of astrophysical particles; hence, constraints can be imposed using only kinematics. On the other hand, electrons and photons of high energy have been detected, and thus the parameter space of the LIVPs of these particles can be restricted from the absence of the vacuum electromagnetic CR and photon decay processes, or the change in the threshold of the photon annihilation process Jacobson_2003 ; Jacobson_2006 .
@@
It should be noted that even though the present work is focused on the study of threshold effects and more specifically particle decays, MDRs have a much broader scope to restrict LIVPs. Constraints in the gravitational sector may be imposed using time of flight delay in multi-messenger detections. At the present only one multi-messenger detection mixing GWs and electromagnetic radiation has been reported, the GW signal GW170817 with the gamma-ray signal from a kilonova Abbott_2017 , but this restrictions using the time delay between GWs and photons have been applied only to impose limits on the Standard Model extension parameters in reference Kosteleck_2015 . Another possibility, which is not considered in the MDR proposed here, is the vacuum birefringence phenomenon caused by the difference in the propagation velocity between the states $h_{+}$ and $h_{\times}$ of GWs.
@@
In conclusion, our work establishes a systematic method to study two body particle decays in a LI violating scenario and impose constraints through high energy astrophysical observations. Extensions to more particle interaction processes can be carried out following the steps exemplified, but these are not of much interest in the gravitational sector. We also want to emphasize the importance of using complementary observations, such as direct detection of GWs, in order to improve the bounds on the parameters introduced. This will be reinforced with future detections of UHECRs and multi-messenger events, which may allow us to understand the elusive nature of QG.
@@ Acknowledgements.
P.M.M. would like to thank Iarley P. Lobo for pointing out to her interesting references. The research of P.M.M. is supported by the project PID2022-138263NB-I00 funded by MICIN/AEI/10.13039/ 501100011033 and by ERDF/EU. The work of J.A.R.C. is partially supported by the project PID2022-139841NB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU, and by COST (European Cooperation in Science and Technology) Actions CA22113 and CA21136. M.A. acknowledges financial support through IPARCOS Master grant.

### v2

@@ 4.2 Observational constraints in the gravitational sector
In order to obtain realistic bounds Kosteleck_2015 , let us consider that the UHECR is a proton of energy $p\sim 10^{11}\,\text{GeV}$. The source of these particles is still unknown Abreu_2022 but it is expected that they are produced in active galactic nuclei; the nearest is found at a few Mpc, so we may take $ct\sim 10\,\text{Mpc}$. Here we do not take into account the spin of the proton and neither its inner structure. Regarding the former, we already have mentioned that spin corrections are not expected to affect the order of magnitude of the constraints on $\mathbb{A}_{\text{proton}}\equiv\mathbb{A}$ and $\mathbb{B}_{\text{graviton}}\equiv\mathbb{B}$. Attending to the latter, a detailed analysis of the gravitational CR emitted by a proton would require us to consider its partonic structure. For soft emitted gravitons with $k\to 0$, this inner structure might be neglected and the pointlike approximation is valid Jacobson_2006 ; for hard emitted gravitons with $k\lesssim p$ it becomes important and the emission rate of its constituents should be computed. This would substantially increase the number of LIVPs considered as $\mathbb{A}_{\text{quark}}\neq\mathbb{A}_{\text{gluon}}$ in general, making the analysis more involved. This study is out of the scope of this work, but it is worth mentioning that this has been studied with more simple MDRs which consider $\alpha=\beta=0$ Gagnon_2004 ; Elliott_2005 . In short, we will assume that the proton has its own effective MDR, whose structure could be understood in terms of the composition of the MDRs of its constituents.
Computing the threshold momentum $k_{\text{th}}$ numerically for $m=m_{\text{proton}}$, performing the integral in Eq. (38) and imposing the condition $\text{d}E/\text{d}t\ll p/t$ we have obtained the allowed $\mathbb{A}-\mathbb{B}$ parameter space for $\alpha=2$, $\beta=0$ and $\alpha=1$, $\beta=2$ shown in Fig. 2. Values of $\mathbb{A}$ and $\mathbb{B}$ that define the dark-grey region permit Cherenkov emission without significant damping, whereas in the light-grey region the process is kinematically allowed but forbidden by the observation because of the damping effect. In both cases we see that no constraints are imposed in the region with $\mathbb{A}>0$ and $\mathbb{B}<0$, and also when $\mathbb{A},\mathbb{B}>0$ and $\mathbb{B}p^{\beta}-\mathbb{A}p^{\alpha}<0$. The latter is due to Eq. (18), since the threshold momentum is $k_{\text{min}}>p$. When both $\mathbb{A}$ and $\mathbb{B}$ are negative the region where kinematics does not permit Cherenkov emission is different in both cases. Attending to Eq. (17), the case where $\alpha=2$ and $\beta=0$ establishes that $\mathbb{B}-3\mathbb{A}p^{2}\geq 0$ so that the threshold momentum is $k_{\text{max}}>0$; by contrast, when $\alpha=1$ and $\beta=2$ the maximum momentum exists when $\mathbb{A}\leq 0$.
@@
The estimated excluded region appears in both cases for $|\mathbb{A}p^{\alpha}|\sim|\mathbb{B}p^{\beta}|\lesssim 10^{-17}$ when $p\sim 10^{11}\,\text{GeV}$ and $ct\sim 10\,\text{Mpc}$. Other positive integer values of $\alpha$ and $\beta$ exhibit constraints of the same order of magnitude for these energies and traveled distances. This allows us to check that LIVPs significantly modify the threshold condition of the process since $|\mathbb{A}p^{\alpha}|\gg(m/p)^{2}$, see Eq. (6). Hence, the approximations performed in Sec. 2 are consistent and the analyses in the massless limit are accurate.
As we have previously advanced, Fig. 2 shows that it is possible to have arbitrarily large values of $\mathbb{A}$ and $\mathbb{B}$ where vacuum gravitational CR is allowed by observation as long as its difference is small enough. We have checked that the same happens for many other positive integer values of $\alpha$ and $\beta$. This fact can be understood in the following way. When computing the energy loss rate, Eq. (38), it is necessary to obtain the integration limits from the threshold condition $\Theta_{c}=0$ for some given $\mathbb{A}$ and $\mathbb{B}$. Considering values of $\mathbb{A}$ and $\mathbb{B}$ where the process is allowed for some $k$, an increase in $\mathbb{A}$ and $\mathbb{B}$ will cause an increase in the value of $\Theta_{c}$ and thus in $\text{d}E/\text{d}t$. However, if the difference between $\mathbb{A}$ and $\mathbb{B}$ is small enough, the region of $k$ where the Cherenkov emission is possible will be narrower; this reduces the value of $\text{d}E/\text{d}t$ as the integration interval is smaller, and it can be such that the condition $\text{d}E/\text{d}t\ll p/t$ is still satisfied. Remember that $\Theta_{c}$ was obtained assuming $|\mathbb{A}p^{\alpha}|\sim|\mathbb{B}k^{\beta}|\ll 1$, and thus our model is not predictive for large values of $\mathbb{A},\mathbb{B}$.
@@ 4.3 Travel distance for particles emitting electromagnetic Cherenkov radiation
We have previously discussed that for positive integer values of $\alpha=\beta$ the energy loss rate of the particle $a$ due to vacuum electromagnetic CR is large enough so that the traveled distance for energies above the threshold $p_{\text{th}}$ is negligible compared to astrophysical distances Coleman_1997 ; Coleman_1999 ; Jacobson_2003 ; Jacobson_2006 . As our MDR does not assume equal values of the exponents of the particle $a$ and the photon ($\alpha$ and $\beta_{\gamma}$, respectively), it could be possible for $a$ to emit CR over large distances. Here we show that, even when $\alpha\neq\beta$, the vacuum electromagnetic CR is much more efficient than its gravitational counterpart, and that the particle $a$ loses all of its energy almost immediately above the threshold.
Figure 3: Constraints on the model for $\alpha=2$, $\beta=0$ (left panel) and $\alpha=1$, $\beta_{\gamma}=2$ (right panel) using an UHECR with $p\sim 10^{11}\,\text{GeV}$ and $ct\sim 0.1\,\text{fm}$. White, light-gray and dark-gray regions represent the same as in Fig. 2 but for the existence or absence of vacuum electromagnetic CR.
To illustrate this, let us consider an ultrahigh energy proton of energy $p\sim 10^{11}\,\text{GeV}$. As we do not solve the threshold Eq. (14) to find the threshold momentum $p_{\text{th}}$ of the particle $a$, we estimate that the order of magnitude of $\mathbb{A}$ and $\mathbb{B}_{\gamma}$ given by the kinematical constraints is the same as the one obtained in the gravitational case; see Fig. 2. For the same values of the exponents $\alpha$ and $\beta_{\gamma}$, Fig. 3 shows the parameter space $\mathbb{A}-\mathbb{B}_{\gamma}$ assuming a traveled distance of $ct\sim 0.1\,\text{fm}$. We see that the region where the damping is not significant lies in distance scales $39$ orders of magnitude smaller compared to those assumed for the emission of vacuum gravitational CR. This means that for UHECRs whose origin is galactic or extragalactic, the emission of electromagnetic CR occurs almost immediately after being produced, losing energy until reaching the threshold momentum $p_{\text{th}}$: the dark-grey region occupies an insignificant amount of the parameter space, and thus the kinematical analysis is enough to obtain bounds on the LIVPs. The estimation was made for two specific pair of values $(\alpha,\beta)$ but we have checked explicitly that similar results are obtained for many other positive integer values.
@@ 5 Conclusions
We would like to remark on an important feature: the parameter space of the LI violating terms is not bounded from above and below, i.e., it is not closed (note that from Fig. 2, $\mathbb{A}$ is bounded from below and $\mathbb{B}$ from above), which implies that gravitational vacuum Cherenkov radiation is permitted as long as the difference between them is small enough. Different processes, such as the decay of a graviton in two high energy particles, could be used to restrict the allowed parameter space, but the quantum nature of gravity has not been resolved yet and no observations of gravitons have been made. Note that this problem is not present in the QED sector. On the one hand, the characteristic time of the analog processes mentioned is much smaller than the traveled distance of astrophysical particles; hence, constraints can be imposed using only kinematics. On the other hand, electrons and photons of high energy have been detected, and thus the parameter space of the LIVPs of these particles can be restricted from the absence of the vacuum electromagnetic CR and photon decay processes, or the change in the threshold of the photon annihilation process Jacobson_2003 ; Jacobson_2006 .
The study carried out along this work is applicable for arbitrary values of the exponents $\alpha$ and $\beta$ until we are interested in deriving additional kinematic features (the characteristics of the solutions to the threshold equation) when both LIVPs $\mathbb{A}$ and $\mathbb{B}$ have equal signs. From this point on we focus on positive values of $\alpha$ and $\beta$, i.e., in violation of LI relevant in the ultraviolet, as expected from many approaches to quantum gravity Addazi_2022 ; batista2023white . Nevertheless, a scenario in which the exponent in the MDR is negative has been studied in Freidel_2021 and Ref.s therein, arguing that quantum gravity phenomenology in the infrared might be possible. In such a case, the phenomenology of UHECR is not relevant for searching of possible signatures of such a scenario. Thus, infrared MDRs do not modify significantly the kinematics of the signatures studied in this work. On the other hand, vacuum gravitational CR associated to low energy processes could serve to impose constraints in the LIVP of the graviton, $\mathbb{B}^{(\beta)}$, when $\beta<0$, as it is possible that the threshold condition admits solutions for arbitrary low momenta depending on the particular MDR under consideration. This interesting study is out of the scope of this work, but we look forward for future projects tackling this aspect.
@@
It should be noted that even though the present work is focused on the study of threshold effects and more specifically particle decays, MDRs have a much broader scope to restrict LIVPs. Constraints in the gravitational sector may be imposed using time of flight delay in multimessenger detections. At the present only one multimessenger detection mixing GWs and electromagnetic radiation has been reported, the GW signal GW170817 with the gamma-ray signal from a kilonova Abbott_2017 , but these restrictions using the time delay between GWs and photons have been applied only to impose limits on the Standard Model extension parameters in Ref. Kosteleck_2015 . Another possibility, which is not considered in the MDR proposed here, is the vacuum birefringence phenomenon caused by the difference in the propagation velocity between the states $h_{+}$ and $h_{\times}$ of GWs.
@@
In conclusion, our work establishes a systematic method to study two body particle decays in a LI violating scenario and impose constraints through high energy astrophysical observations. Extensions to more particle interaction processes can be carried out following the steps exemplified, but these are not of much interest in the gravitational sector. We also want to emphasize the importance of using complementary observations, such as direct detection of GWs, in order to improve the bounds on the parameters introduced. This will be reinforced with future detections of UHECRs and multimessenger events, which may allow us to understand the elusive nature of QG.
@@ Acknowledgements.
P.M.M. would like to thank Iarley P. Lobo for pointing out to her interesting Ref.s. The research of P.M.M. is supported by the project PID2022-138263NB-I00 funded by MICIN/AEI/10.13039/ 501100011033 and by ERDF/EU. The work of J.A.R.C. is partially supported by the project PID2022-139841NB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU, and by COST (European Cooperation in Science and Technology) Actions CA22113 and CA23130. M.A. acknowledges financial support through IPARCOS Master grant.